Two forces, 3 N and 5 N respectively, act on the same point simultaneously. The magnitude of their resultant force is 2 N when the angle between them is _____.


1
Expert's answer
2021-11-29T11:45:44-0500

The resultant of two vectors:

"\vec{P}" and "\vec{Q}" is given by

"R= \sqrt{ P^2+Q^2+2PQcosθ}"

where, P and Q are magnitudes of two vectors and θ is the angle between two vectors.

Here θ=? P=3 N Q=5 N R=2 N

Therefore,

"2 = \sqrt{3^2 + 5^2 +2 \times 3 \times 5 \times cosθ} \\

4 = 9 + 25 + 30cosθ \\

30cosθ = -30 \\

cosθ = -1 \\

θ = 180°"